OpenStudy (anonymous):

The derivative of f(x,y) at P(1,2) in the direction of i+j is 2sqrt(2) and in the direction of -2j is -3. What is the derivative of f in the direction of -i-2j?

OpenStudy (anonymous):

Suppose the gradient of f is (x1,y1) at (1,2) then (x1,y1) . (1,1)/sqrt(2) = 2 sqrt(2) Hence x1+y1 = 4 from the second directional derivative, we get -2 y1 =-6 Hence the gradient of f at (1,2) is (1,3) Now compute (1,3) . (-1,-2) and divide sqrt(5) to get \[-\frac{7}{\sqrt{5}} \]

OpenStudy (anonymous):

We used that the directional derivative in the direction of a vector u is the dot product of the gradient of at the point and a unit vectr v in the direction of u.

OpenStudy (anonymous):

thank you